# Equilibrium Investment with Random Risk Aversion: (Non-)uniqueness, Optimality, and Comparative Statics

### v2

@@ Abstract
This paper investigates infinite-dimensional portfolio selection problem under a general distribution of the risk aversion parameter. We provide a complete characterization of all deterministic equilibrium investment strategies. Our results reveal that the solution structure depends critically on the distribution of risk aversion: the equilibrium is unique whenever it exists in the case of finite expected risk aversion, whereas an infinite expectation can lead to infinitely many equilibria or to a unique trivial one ($\bar{\pi}\equiv\mathbf{0}$). To address this multiplicity, we introduce three optimality criteria—optimal, uniformly optimal, and uniformly strictly optimal—and explicitly characterize the existence and uniqueness of the corresponding equilibria. Under the same necessary and sufficient condition, the optimal and uniformly optimal equilibria exist uniquely and coincide. Furthermore, assuming that the market price of risk is non-zero near the terminal time, we show that the optimal (and hence uniformly optimal) equilibrium is also uniformly strictly optimal. Finally, we perform comparative statics to demonstrate that a risk aversion distribution dominating another in the reverse hazard rate order leads to a less aggressive equilibrium strategy.
@@ 1 Introduction
The classical framework for portfolio selection, pioneered by merton1969lifetime; merton1971 and samuelson1969lifetime, is based on the assumption that an investor’s risk aversion can be captured by a single known constant. In practice, however, this coefficient is very difficult for an agent to specify with certainty, and a growing body of empirical evidence suggests that it varies dynamically with market conditions, often increasing during financial crises and bear markets, as shown in studies by gordon2000preference; chetty2006new and guiso2018time. These challenges have motivated a move towards models that incorporate random risk aversion. Within this paradigm, a more economically coherent approach, as advanced by Desmettre2023, is to aggregate certainty equivalents over the distribution of the risk aversion parameter. While the random risk aversion in Desmettre2023 provides a consistent performance measure in monetary units, it introduces a nonlinearity in expectation that naturally leads to the time inconsistency of the preference.
@@
The first systematic treatment of time inconsistency was given by Strotz1955, who proposed the notion of consistent planning, laying the foundation for the intra-personal game-theoretic framework. Building on this idea, ekeland2006being later developed a rigorous continuous-time formalization, which in turn inspired extensive research on continuous-time control with time-inconsistent preferences. For example, bjork2010general (published version Bjork2017) established a general theoretical framework by deriving an extended HJB equation and hu2012time; Hu2017 investigated a time-inconsistent stochastic linear–quadratic control problem. For further discussions and related developments, see, e.g., yan2019time, he2021equilibrium, Hernandez2023, and the references therein.
@@
Within the Black-Scholes market model, Desmettre2023 provide a verification theorem on the intra-personal equilibrium strategies for a CRRA utility function with a general random risk aversion and, in the case of a binomial random risk aversion, characterize the equilibrium by a three-dimensional ODE system without establishing the existence and uniqueness of the solution. [^1] This work has spurred further research extending the model to more complex settings. For example, chen2025equilibrium incorporate a regime-switching framework in which market dynamics and preferences co-evolve; WANG2025103140 study the management of the DC pension plan, providing semi-explicit solutions for a CRRA utility with a binomial random risk aversion; he2025dynamic establish the existence and uniqueness of the solution to an infinite-dimensional Riccati equation which characterizes the equilibrium strategies under Heston’s SV model for a CRRA or CARA utility function with a general bounded random risk aversion. For an investigation on the corresponding pre-commitment problem, see, e.g., xia2024optimal. This technical gap is filled by liang2025short. For general random risk aversion, liang2025short characterize intra-personal equilibrium strategies using an integral equation. They establish the existence and uniqueness of the solution to the integral equation under the assumptions that the variance of the random risk aversion is finite and that the function $h$ defined in (2.6) below is bounded.
@@
This paper aims to offer more insight into the portfolio selection problem with random risk aversion within the simplest model, in line with Desmettre2023 and liang2025short. Our main contributions are summarized as follows:
@@
  • Complete characterization of all deterministic equilibria. We relax the assumptions in the existing literature and derive a full characterization of all deterministic equilibrium strategies. An important finding is that there may be infinitely many equilibria when the expectation of the random risk aversion is infinite. The uniqueness of the equilibrium strategy for the continuous-time time-inconsistent control problem is a very challenging problem (see, e.g., hjz2021). This paper sheds light on this issue by demonstrating that equilibrium strategies, even within the class of deterministic strategies, may lack uniqueness. Furthermore, we provide a necessary and sufficient condition for the uniqueness of the deterministic equilibrium.
@@
  • Resolution of equilibrium multiplicity via the optimality criterion. A question then arises: which one should the agent choose from these infinitely many equilibrium strategies? We suggest that the agent should maximize the objective functional at the initial time over all available equilibria. This leads to the discussion of the so-called optimal equilibria. To our best knowledge, there is only one existing paper, wei2024time, on optimal equilibria for a time-inconsistent control problem, and there are three papers, hz2019; hz2020mf and hw2021, on optimal equilibria for time-inconsistent stopping problems. We give a necessary and sufficient condition for the existence of optimal equilibria and observe that the optimal equilibrium, if exists, is unique and is also uniformly optimal. Moreover, if the market price of risk remains non-zero near the terminal time, the optimal equilibrium is also uniformly strictly optimal.
@@
  • Comparative statics on equilibrium strategies. We examine how the distribution of the random risk aversion influences the equilibrium strategies. It is natural to conjecture that a “larger” random risk aversion would lead to less risky investments. We observe by counterexamples that the “largeness” in the sense of the first-order stochastic dominance is insufficient to validate the conjecture. However, a stronger stochastic order, the reverse hazard rate dominance, is sufficient. Thus we get an analogy with Borell2007 and xia2011risk, where the comparative statics on the optimal portfolio with respect to risk aversion are carried out for the standard expected utility maximization with a general utility function and within the Black-Scholes market model (The corresponding comparative statics for a static model with a risk-free asset and a risky asset goes back to Arrow1970 and Pratt1964:RiskAversion).
The paper is organized as follows. Section 2 presents the model and derives the integral equation characterizing the equilibrium. Section 3 analyzes both the finite and infinite expectation cases, providing a complete characterization of all equilibria. Section 4 introduces three optimality criteria for equilibrium selection and establishes the necessary and sufficient conditions for their existence and uniqueness. Section 5 performs comparative statics to investigate the impact of risk aversion distributions on equilibrium strategies, establishing monotonicity under the reverse hazard rate order, and provides numerical illustrations that verify the theoretical results and present a counterexample under first-order stochastic dominance. [^2] Herein, “increasing” means “non-decreasing” and “decreasing” means “non-increasing.”
@@ 2 Problem Formulation
Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t},\ 0\leq t\leq T\}$ is the augmented natural filtration generated by a $d$-dimensional standard Brownian motion $\{B(t),\ 0\leq t\leq T\}$ and $\mathcal{F}=\mathcal{F}_{T}$. The market consists of one risk-free asset (bond) and $d$ risky assets (stocks). For simplicity, we assume that the interest rate of the bond is zero. The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$, are given by
@@
and $\sigma(t)$ is invertible for every $t\in[0,T]$. The market price of risk is $\lambda(t)\triangleq(\sigma(t))^{-1}\mu(t)$, $t\in[0,T]$. We also assume that $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t<\infty$.
For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes. For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables.
A trading strategy is a process $\pi=\{\pi_{t},\ t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that
@@
where $\pi_{t}$ represents the vector of portfolio weights determining the investment of wealth in the stocks at time $t$. For any $t\in[0,T)$ and any initial wealth level $w>0$, let $W^{t,w,\pi}=\{W^{t,w,\pi}_{s},\ t\leq s\leq T\}$ denote the self-financing wealth process starting from $w$ at time $t$ and driven by a trading strategy $\pi$. It evolves according to the stochastic differential equation (SDE):
$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\pi}_{s}=W^{t,w,\pi}_{s}\pi^{\top}_{s}\mu(s)\mathrm{d}s+W^{t,w,\pi}_{s}\pi^{\top}_{s}\sigma(s)\mathrm{d}B(s),\quad s\geq t,\\ &W^{t,w,\pi}_{t}=w.\end{aligned}\right$. (2.1)
@@
$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},&\gamma>0,\gamma\neq 1,x>0,\\ \log x,&\gamma=1,x>0.\end{cases}$
Under the expected utility theory, for each CRRA utility function $u^{\gamma}$, at time $t$, the certainty equivalent of $W^{t,w,\pi}_{T}$ is $\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)$. Following Desmettre2023, we assume that the objective functional of the agent at time $t$ is the weighted average of the certainty equivalents:
@@
$\begin{cases}\mathbb{E}_{t}\left[\left|u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right|\right]<\infty\,\ a.s.\quad\text{ for all }\gamma\in\mathrm{supp}(F_{\bm{R}}),\\ \int_{[0,\infty)}\left|\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\right|\mathrm{d}F_{\bm{R}}(\gamma)<\infty\,\ \ a.s..\end{cases}$
@@
The strategy $\pi^{t,\varepsilon,k}$ serves as a perturbation of $\bar{\pi}$. Following hu2012time; Hu2017 and yan2019time, we introduce the definition of equilibrium strategies as follows.
@@
A strategy $\bar{\pi}$ is called an equilibrium strategy if for any $t\in[0,T)$, $w>0$, and $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, whenever $\pi^{t,\varepsilon,k}\in\Pi$ for all sufficiently small $\varepsilon>0$, and for any positive sequence $\{\varepsilon_{n},n\geq 1\}$ satisfying $\lim\limits_{n\to\infty}\varepsilon_{n}=0$ we have
$\limsup_{n\to\infty}\frac{J(\pi^{t,\varepsilon_{n},k};t,w)-J(\bar{\pi};t,w)}{\varepsilon_{n}}\leq 0\quad\text{a.s.}.$ (2.3)
@@
In literature, the inequality is usually
$\limsup_{\varepsilon\downarrow 0}\frac{J(\pi^{t,\varepsilon,k};t,w)-J(\bar{\pi};t,w)}{\varepsilon}\leq 0\quad\text{a.s.}.$
However, the left-hand side of the inequality might be unmeasurable. To guarantee the measurability, we modify the definition of equilibrium accordingly.
@@
$J_{0}(t,\pi)\triangleq\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\tilde{\mathbb{E}}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}/W^{\pi}_{t}}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma)$
is the objective functional used in liang2025short and $W^{\pi}$ evolves according to
$\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\pi^{\top}_{t}\mu(t)\mathrm{d}t+W^{\pi}_{t}\pi^{\top}_{t}\sigma(t)\mathrm{d}B(t),\quad W^{\pi}_{0}=w_{0}>0.$
@@
In our analysis, we follow the approach of liang2025short and focus on the equilibrium strategies given by $\bar{\pi}=(\sigma^{\top})^{-1}a$, where $a$, referred to as the risk exposure vector, is a deterministic and right-continuous funcion in the $L^{2}$ space. Let $\Pi_{d}\subset\Pi$ denote the set of all such strategies. Because $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, the perturbed strategy $\pi^{t,\varepsilon,k}$ can be random, even though $\bar{\pi}\in\Pi_{d}$ itself is deterministic. The wealth process $W^{t,w,\bar{\pi}}$ of the portfolio $\bar{\pi}$ satisfies the following SDE
$\mathrm{d}W^{t,w,\bar{\pi}}_{s}=W^{t,w,\bar{\pi}}_{s}[a(s)^{\top}\lambda(s)\mathrm{d}s+a(s)^{\top}\mathrm{d}B(s)],\quad s\geq t,$
@@
The following theorem shows that $\pi\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$. (This result is natural: if the random risk aversion is very large, then the investor puts no money into the risky asset. However, in Theorem 3.2(2) we will see that $\pi\equiv\mathbf{0}$ is not necessarily the unique equilibrium.)
@@
$\bar{\pi}\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
@@
$=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma),$
@@
$=\left(\mu^{\top}(t)k-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma)\right).$
@@
$\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$ is an equilibrium if and only if
@@
holds for any $t\in[0,T)$, where [^3] This function $h$ is slightly different from that in liang2025short: $h(x)$ there amounts to $h(x^{2})$ here. Nevertheless, the relevant properties, such as the boundedness of $h$, remain equivalent under the two formulations.
@@
When $\tilde{\mathbb{E}}[\bm{R}]<\infty$, we have $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}\in(0,\infty)$ and $h\in C([0,\infty))$. In this case, an argument similar to the proof of liang2025short yields the equivalence.
When $\tilde{\mathbb{E}}[\bm{R}]=\infty$, we have $h(0)=0$. In this case, similar to the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, (2.3) and (2.5) are equivalent for $t\in[0,T_{0})$, where $T_{0}=\inf\{t\in[0,T]:v_{a}(t)=0\}$. Moreover, they are automatically satisfied for $t\in[T_{0},T)$ based on Theorem 2.4 and $h(0)=0$. ∎
@@
The function $h$ defined in (2.6) plays a central role in our analysis. Here, we present two examples for which $h$ can be explicitly worked out.
@@
Consider the following two distributions for the random risk aversion $\bm{R}$.
@@
In the two cases of the above example, the functions $h$ are unbounded, violating the boundedness assumption on $h$ made by liang2025short. Moreover, as we will see in Example 3.5 below, the equilibrium strategies can be explicitly obtained for both cases. To accommodate such interesting cases, we will conduct a thorough analysis on Condition (2.5) and the equilibrium strategies under a general distribution of $\bm{R}$.

### v3

@@ Abstract
This paper studies a continuous-time portfolio selection problem under a general distribution of random risk aversion (RRA). We provide a complete characterization of all deterministic equilibrium strategies in closed form. Our results show that the structure of the solution depends crucially on the distribution of RRA: the equilibrium is unique (if exits) when the expectation of RRA is finite, whereas an infinite expectation leads either to infinitely many equilibria or to a unique trivial one (i.e. risk-free investment). To resolve this multiplicity of equilibria, we select, among all deterministic equilibria, the one that maximizes the objective functional at the initial time. We establish a necessary and sufficient condition for the existence of such an optimal equilibrium, which is then shown to be unique and uniformly optimal. Finally, we conduct a comparative statics. Using counterexamples based on two-point distributed RRA, we demonstrate that a larger risk aversion in the sense of first-order stochastic dominance does not necessarily lead to less risky investment. Within the two-point distribution framework, we further examine the single-crossing property of equilibrium strategies and the monotonicity of the crossing time. We show that a larger risk aversion under a stronger stochastic order—the reverse hazard rate order—always leads to less risky investment. In addition, we analyze how the convex combination of independent and identically distributed RRAs influences investment.
@@ 1 Introduction
The classical framework for portfolio selection, pioneered by Merton (1969, 1971) and Samuelson (1969), is based on the assumption that an investor’s risk aversion can be captured by a single known constant. In practice, however, this coefficient is very difficult for an agent to specify with certainty, and a growing body of empirical evidence suggests that it varies dynamically with market conditions, often increasing during financial crises and bear markets, as shown in studies by Gordon and St-Amour (2000), Chetty (2006), and Guiso et al. (2018). These challenges have motivated a move towards models that incorporate random risk aversion (RRA). Within this paradigm, a more economically coherent approach, as advanced by Desmettre and Steffensen (2023), is to aggregate the certainty equivalents over the distribution of the risk aversion parameter. While the aggregation in Desmettre and Steffensen (2023) provides a consistent performance measure in monetary units, it introduces a nonlinearity in expectation, which naturally leads to the time inconsistency of the preference.
@@
The first systematic treatment of time inconsistency was given by Strotz (1955), who proposed the notion of consistent planning, laying the foundation for the intra-personal game-theoretic framework. Building on this idea, Ekeland and Lazrak (2006) later developed a rigorous continuous-time formalization, which in turn inspired extensive research on continuous-time control with time-inconsistent preferences. For example, Bjork and Murgoci (2010) (published version Björk et al. (2017)) established a general theoretical framework by deriving an extended HJB equation and Hu, Jin, and Zhou (2012, 2017) investigated a time-inconsistent stochastic linear–quadratic control problem. For further discussions and related developments, see, e.g., Yan and Yong (2019), He and Jiang (2021), Hernández and Possamaï (2023), and the references therein.
@@
Within the Black-Scholes market model, Desmettre and Steffensen (2023) provide a verification theorem on the intra-personal equilibrium strategies for a general RRA and, in the case of a two-point distributed RRA, characterize the equilibrium by a three-dimensional ODE system without establishing the existence and uniqueness of the solution. [^1] This work has spurred further research extending the model to more complex settings. For example, Chen, Guan, and Liang (2025) incorporate a regime-switching framework in which market dynamics and preferences co-evolve; Aquino et al. (2025) study a continuous-time portfolio choice problem where state-dependent preferences are determined by an exogenous factor evolving as an Itô diffusion; Wang and Jia (2025) study the management of the DC pension plan, providing semi-explicit solutions for a two-point distributed RRA; He, Jiang, and Xia (2025) establish the existence and uniqueness of the solution to an infinite-dimensional Riccati equation which characterizes the equilibrium strategies under Heston’s SV model for a general bounded RRA. For an investigation on the corresponding pre-commitment problem, see, e.g., Xia (2024). This technical gap is filled by Liang, Wang, and Xia (2025a). For a general RRA, Liang, Wang, and Xia (2025a) characterize intra-personal equilibrium strategies using an integral equation. They establish the existence and uniqueness of the solution to the integral equation under the assumptions that the variance of the RRA is finite and that the function $h$ defined in (2.6) below is bounded.
@@
This paper aims to provide deeper insight into the portfolio selection problem with RRA within the simplest modeling framework, in line with Desmettre and Steffensen (2023) and Liang, Wang, and Xia (2025a). Our main contributions are summarized below.
@@
  • Complete characterization of all deterministic equilibria in closed form. We relax the assumptions in the existing literature and derive, in closed form, all deterministic equilibrium strategies. An important finding is that there may be infinitely many equilibria when the expectation of the RRA is infinite. The uniqueness of the equilibrium strategy for the continuous-time time-inconsistent control problem is a very challenging problem (see, e.g., Hu, Jin, and Zhou (2021, Section 7)). This paper sheds light on this issue by demonstrating that equilibrium strategies, even within the class of deterministic strategies, may lack uniqueness. Furthermore, we provide a necessary and sufficient condition for the uniqueness of the deterministic equilibrium.
@@
  • Resolution of equilibrium multiplicity via the optimality criterion. A question then arises: which one should the agent choose from these infinitely many equilibrium strategies? We suggest that the agent should maximize the objective functional at the initial time over all available equilibria. This leads to a discussion of the so-called optimal equilibria. To our best knowledge, there is only one existing paper, Wei, Xia, and Zhao (2026), on optimal equilibria for a time-inconsistent control problem, and there are three papers, Huang and Zhou (2019, 2020) and Huang and Wang (2021), on optimal equilibria for time-inconsistent stopping problems. We give a necessary and sufficient condition for the existence of optimal equilibria and observe that the optimal equilibrium, if it exists, is unique and is also uniformly optimal. Moreover, if the market price of risk remains non-zero near the terminal time, the optimal equilibrium is also uniformly strictly optimal.
@@
  • Comparative statics on equilibrium strategies. We investigate how the distribution of RRA affects the equilibrium strategies. Intuitively, one would expect a “comparative property”: a “larger” RRA should lead to less risky investment. However, through counterexamples involving two-point distributed RRA, we demonstrate that “largeness” in the sense of first-order stochastic dominance is insufficient to guarantee this property. Within the same two-point distribution framework, we further examine the single-crossing property of equilibrium strategies and the monotonicity of the crossing time with respect to (w.r.t.) the probability parameter. In contrast, a stronger stochastic order—reverse hazard rate dominance—is shown to be sufficient for the comparative property to hold. This yields an analogy with Borell (2007) and Xia (2011), who conducted comparative statics on optimal portfolios w.r.t. risk aversion in the context of expected utility maximization with general utility functions under the Black–Scholes market model. (The corresponding analysis for a static model with one risk-free asset and one risky asset dates back to Arrow (1970) and Pratt (1964).) Finally, we show that a convex combination of independent and identically distributed (i.i.d.) RRAs results in less risky investment.
The rest of the paper is organized as follows. Section 2 introduces the model and derives the integral equation that characterizes the equilibrium. Section 3 analyzes both finite- and infinite-expectation cases, offering closed-form solutions for all equilibria. Section 4 resolves equilibrium multiplicity by proposing an optimality-based selection criterion, and establishes necessary and sufficient conditions for the existence, uniqueness, and uniform (strict) optimality of optimal equilibria. Section 5 performs a comparative statics of equilibrium strategies w.r.t. the distribution and the aggregation of RRA, identifying conditions under which “larger” risk aversion leads to less risky investment.
@@ 2 Problem Formulation
Let $T>0$ be a finite time horizon and $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$ a filtered complete probability space, where $\mathbb{F}=\{\mathcal{F}_{t}:\ 0\leq t\leq T\}$ is the augmented natural filtration generated by a $d$-dimensional standard Brownian motion $\{B(t):\ 0\leq t\leq T\}$ and $\mathcal{F}=\mathcal{F}_{T}$. The market consists of one risk-free asset (bond) and $d$ risky assets (stocks). For simplicity, we assume that the interest rate of the bond is zero. The dynamics of the stock price processes $S_{i}$, $i=1,\cdots,d$, is given by
@@
and that $\sigma(t)$ is invertible for every $t\in[0,T]$. The market price of risk is $\lambda(t)\triangleq(\sigma(t))^{-1}\mu(t)$, $t\in[0,T]$. We also assume that $\int_{0}^{T}|\lambda(t)|^{2}\mathrm{d}t<\infty$.
For any $m\geq 1$ and $\mathbb{S}\subset\mathbb{R}^{m}$, $L^{0}(\mathbb{F},\mathbb{S})$ is the space consisting of $\mathbb{S}$-valued, $\mathbb{F}$-progressively measurable processes. For each $t\in[0,T]$, $p\in[1,\infty]$, $L^{p}(\mathcal{F}_{t},\mathbb{S})$ is the set of all $L^{p}$-integrable, $\mathbb{S}$-valued, and $\mathcal{F}_{t}$-measurable random variables.
A trading strategy is a process $\pi=\{\pi_{t}:\ t\in[0,T)\}\in L^{0}(\mathbb{F},\mathbb{R}^{d})$ such that
@@
where $\pi_{t}$ is the vector of portfolio weights at time $t$. For any $t\in[0,T)$ and any initial wealth level $w>0$, let $W^{t,w,\pi}=\{W^{t,w,\pi}_{s}:\ t\leq s\leq T\}$ denote the self-financing wealth process starting from $w$ at time $t$ and driven by a trading strategy $\pi$. It evolves according to the stochastic differential equation (SDE):
$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\pi}_{s}=W^{t,w,\pi}_{s}\pi^{\top}_{s}\mu(s)\mathrm{d}s+W^{t,w,\pi}_{s}\pi^{\top}_{s}\sigma(s)\mathrm{d}B(s),\quad s\in[t,T],\\ &W^{t,w,\pi}_{t}=w.\end{aligned}\right$. (2.1)
@@
$u^{\gamma}(x)=\begin{cases}\frac{x^{1-\gamma}}{1-\gamma},&\gamma\geq 0,\gamma\neq 1,x>0,\\ \log x,&\gamma=1,x>0.\end{cases}$
Under the expected utility theory, for each CRRA utility function $u^{\gamma}$, at time $t$, the certainty equivalent of $W^{t,w,\pi}_{T}$ is $\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)$, where $\mathbb{E}_{t}$ denotes the conditional expectation w.r.t. $\mathcal{F}_{t}$. Following Desmettre and Steffensen (2023), we assume that the objective functional of the agent at time $t$ is the weighted average of the certainty equivalents:
@@
$\begin{cases}\mathbb{E}_{t}\left[\left|u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right|\right]<\infty\,\ a.s.\quad\text{ for all }\gamma\in\operatorname*{supp}(F_{\bm{R}}),\\ \displaystyle\int_{[0,\infty)}\left|\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left(W^{t,w,\pi}_{T}\right)\right]\right)\right|\mathrm{d}F_{\bm{R}}(\gamma)<\infty\,\ \ a.s.\end{cases}$
@@
The strategy $\pi^{t,\varepsilon,k}$ serves as a perturbation of $\bar{\pi}$. Following Hu, Jin, and Zhou (2012, 2017) and Yan and Yong (2019), we introduce the definition of equilibrium strategies as follows.
@@
A strategy $\bar{\pi}$ is called an equilibrium strategy if, for any $t\in[0,T)$, $w>0$, $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$ with $\pi^{t,\varepsilon,k}\in\Pi$ for all sufficiently small $\varepsilon>0$, and any positive sequence $\{\varepsilon_{n},n\geq 1\}$ with $\lim\limits_{n\to\infty}\varepsilon_{n}=0$, we have
$\limsup_{n\to\infty}\frac{J(\pi^{t,\varepsilon_{n},k};t,w)-J(\bar{\pi};t,w)}{\varepsilon_{n}}\leq 0\quad\text{a.s.}$ (2.3)
@@
In literature, the following inequality
$\limsup_{\varepsilon\downarrow 0}\frac{J(\pi^{t,\varepsilon,k};t,w)-J(\bar{\pi};t,w)}{\varepsilon}\leq 0\quad\text{a.s.}$
is usually used to define equilibrium strategies. However, the left-hand side of the inequality might be unmeasurable. To guarantee the measurability, we use the upper limit of a sequence instead of the upper limit of an uncountable net.
@@
$J_{0}(t,\pi)\triangleq\int_{[0,\infty)}\left(u^{\gamma}\right)^{-1}\left(\mathbb{E}_{t}\left[u^{\gamma}\left({W^{\pi}_{T}/W^{\pi}_{t}}\right)\right]\right)\mathrm{d}F_{\bm{R}}(\gamma)$
is the objective functional used in Liang, Wang, and Xia (2025a) and $W^{\pi}$ evolves according to
$\left\{\begin{aligned} &\mathrm{d}W^{\pi}_{t}=W^{\pi}_{t}\pi^{\top}_{t}\mu(t)\mathrm{d}t+W^{\pi}_{t}\pi^{\top}_{t}\sigma(t)\mathrm{d}B(t),\quad t\in[0,T],\\ &W^{\pi}_{0}=w_{0}>0.\end{aligned}\right$.
@@
In our analysis, we follow the approach of Liang, Wang, and Xia (2025a) and focus on the equilibrium strategies given by $\bar{\pi}=(\sigma^{\top})^{-1}a$, where $a$, referred to as the risk exposure vector, is a deterministic and right-continuous function in the $L^{2}$-space. Let $\Pi_{d}\subset\Pi$ denote the set of all such strategies. Because $k\in L^{\infty}(\mathcal{F}_{t},\mathbb{R}^{d})$, the perturbed strategy $\pi^{t,\varepsilon,k}$ can be random, even though $\bar{\pi}\in\Pi_{d}$ itself is deterministic. The wealth process $W^{t,w,\bar{\pi}}$ of the portfolio $\bar{\pi}$ satisfies the following SDE
$\left\{\begin{aligned} &\mathrm{d}W^{t,w,\bar{\pi}}_{s}=W^{t,w,\bar{\pi}}_{s}[a(s)^{\top}\lambda(s)\mathrm{d}s+a(s)^{\top}\mathrm{d}B(s)],\quad s\in[t,T],\\ &W^{t,w,\bar{\pi}}_{t}=w,\end{aligned}\right$.
@@
The following theorem shows that $\bar{\pi}\equiv\mathbf{0}$ is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$. (This result is natural: if the RRA is very large, then the investor puts no money into the risky asset. However, in Theorem 3.2(2) we will see that $\bar{\pi}\equiv\mathbf{0}$ is not necessarily the unique equilibrium.)
@@
The risk-free investment ($\bar{\pi}\equiv\mathbf{0}$) is an equilibrium if $\tilde{\mathbb{E}}[\bm{R}]=\infty$.
@@
$=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\int_{[0,\infty)}\gamma\mathrm{d}F_{\bm{R}}(\gamma)=-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\tilde{\mathbb{E}}[\bm{R}],$
@@
$=\left(\mu^{\top}(t)k-\frac{1}{2}|\sigma^{\top}(t)k|^{2}\tilde{\mathbb{E}}[\bm{R}]\right).$
@@
A strategy $\bar{\pi}=(\sigma^{\top})^{-1}a\in\Pi_{d}$ is an equilibrium if and only if
@@
holds for any $t\in[0,T)$, where [^3] This function $h$ is slightly different from that one in Liang, Wang, and Xia (2025a): $h(x)$ there amounts to $h(x^{2})$ here. Nevertheless, the relevant properties, such as the boundedness of $h$, remain equivalent under the two formulations.
@@
When $\tilde{\mathbb{E}}[\bm{R}]<\infty$, we have $h(0)={1}/{\tilde{\mathbb{E}}[\bm{R}]}\in(0,\infty)$ and $h\in C([0,\infty))$. In this case, an argument similar to the proof of Liang, Wang, and Xia (2025a, Theorem 3.4) yields the equivalence.
When $\tilde{\mathbb{E}}[\bm{R}]=\infty$, we have $h(0)=0$. In this case, similar to the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, (2.3) and (2.5) are equivalent for $t\in[0,T_{0})$, where $T_{0}\triangleq\inf\{t\in[0,T]:\ v_{a}(t)=0\}$. Moreover, they are automatically satisfied for $t\in[T_{0},T)$ based on Theorem 2.4 and $h(0)=0$. ∎
@@
Its derivative $l^{\prime}(0)$ exists and is finite for all $y>0$. In the case $\tilde{\mathbb{E}}[\bm{R}]<\infty$, we have $l^{\prime}(0)=-\tilde{\mathbb{E}}[\bm{R}]$. In the case $\tilde{\mathbb{E}}[\bm{R}]=\infty$, we extend the definition by setting $l^{\prime}(0)=$-$\infty$. Then, for any $x\geq 0$, the function $h$ is related to $l$ through $h(x)=-{l(x/2)}/{l^{\prime}(x/2)}$. The function $h$ defined in (2.6) plays a central role in our analysis. Here, we present two examples for which $h$ can be explicitly worked out.
@@
Consider the following two distributions for the RRA $\bm{R}$.
@@
In the two cases of the above example, the functions $h$ are unbounded, violating the boundedness assumption on $h$ made by Liang, Wang, and Xia (2025a). Moreover, as we will see in Example 3.5 below, the equilibrium strategies can be explicitly obtained for both cases. To accommodate such interesting cases, we will conduct a thorough analysis on Condition (2.5) and the equilibrium strategies under a general distribution of $\bm{R}$.